# Engagement Maximization

### v5

@@ 2.3 Discussion
The generalized entropy functions $G$ and $H$ play conceptually distinct roles in our model. The function $H$ governs the agent’s information processing constraint. It determines which kinds of information are relatively easy or difficult for the agent to process. The function $G$, on the other hand, measures the kind of engagement that generates profits for the principal. One leading case of interest is when $H=G$.
In the context of an internet platform, we interpret this case as one in which the principal displays advertisements to the user interspersed with content, and the principal is paid “per impression” by advertisers without any direct measurement of whether the user paid attention to the advertisement. The principal’s profits are maximized by maximizing the time the agent spends attending to the principal’s content. Arrangements similar to this assumption are observed in a variety of media markets (see, e.g., Gentzkow et al. (2024)). [^10] The authors would like to thank Emir Kamenica for suggesting this interpretation. Note that when $H=G$, the principal’s objective is not to maximize the stopping time $\tau$ directly; however, in equilibrium the constraint (2) will bind, and the principal will maximize time spent spend subject to (2) holding with equality. In the context of a teacher attempting to engage a test-motivated student, we interpret the $H=G$ case as one in which the teacher’s goal is for the student to learn as much as possible, without any particular preference for whether the student learns test-relevant or test-irrelevant information.
A second leading case is one in which $G$ measures only information acquisition that is irrelevant to the agent’s decision problem. In the context of an internet platform, we interpret this case as one in which the platform profits from the users’ attention to advertisements that are unrelated to the content the user is interested in. This case differs from the $H=G$ case in that the internet platform profits only if the user actually pays attention to the advertisement. This would be the case if there is some measurement of the effectiveness of the ad (using clicks, cookies, or other trackers), and the platform earns revenue on the basis of this measurement. In the context of engaging test-motivated students, we interpret this case as one in which the teacher cares only about whether the student learns the course material and not about whether the student does well on the test, which is the exact opposite of what the student cares about.
In both the internet platform and teacher-student context, it is natural to assume that elements of both of these two leading cases are present. An internet platform will profit on both a per-impression basis and on the basis of whether ads on its platform are effective, which can be modeled in our framework as setting $G$ equal to the sum of $H$ and another entropy function measuring the attention paid to advertisements. This second component of $G$ could be irrelevant to the agent’s decision problem, if the ads are in fact unrelated, or not, if the ads are related to information the user is attempting to acquire. Likewise, in the teacher-student context, a teacher might care both about whether the student passes the test and whether the student learns the course material (including content that can and cannot be tested in an exam). This can be modeled in our framework as setting $G$ equal to the sum of $\hat{u}$ (representing the student’s expected test performance) and a function measuring the student’s understanding of the course content. [^11] Our preferred interpretation of the model is that the principal cares only about the information the agent acquires and not the action the agent selects. However, this example shows that our model also covers the case in which the principal and agent agree on the optimal action conditional on the agent’s beliefs (although our extension to the case of an informed principal in Section 5.2 would not hold in this case). Our assumption that $A$ is finite implies that $\hat{u}$ is not continuously differentiable, and hence this example would require minor modifications of some of our technical assumptions.
Apart from the distinction between $H$ and $G$, our formulation of the principal’s problem (Definition 1) embeds several assumptions. First, we assume that the principal acts first and can commit to law $P$ before the agent chooses her stopping strategy $\tau$. Given that the agent faces a pre-determined $P$, her problem is dynamically consistent and the agent’s commitment or lack thereof is irrelevant. In the context of an internet platform and a user, this assumption can be motivated by assuming that the internet platform designs its recommendation algorithms before the user can commit to a usage pattern. However, this justification is inconsistent with the platform’s ability to change its algorithms at any time without informing its users, and is not particularly appealing in the teacher-student context either. For this reason, we will devote a significant part of our analysis to the case without commitment in Section 4.2, and demonstrate that the solutions we construct in the commitment case can also be used to characterize equilibria in the case without commitment.
Second, we assume that the belief process is a martingale under both the principal and the agent’s initial information set ($\mathcal{F}_{0}$). This assumption is standard in the literature on Bayesian Persuasion (following Kamenica and Gentzkow (2011)) and is usually interpreted as assuming that the principal commits to a particular information design before learning the true state. In the typical Bayesian persuasion setting, this assumption is important; a sender whose preferences are ex-post aligned with the receiver would like to truthfully reveal the state, and might be able to do so by signaling via the choice of information structure. As a result, the lack of commitment to the information structure strictly restricts what the sender can implement (see Lipnowski et al. (2022); Min (2021)). In contrast, in our model, an informed principal would never want to reveal the true state, as doing so would cause the agent to immediately stop, and for this reason our main results apply to the case of an informed principal, as discussed in Section 5.2.
Third, we assume that the agent attends fully to the information the principal provides (i.e. that the $P$ measure chosen by the principal is that one that generates the agent’s expectations). We show in Section 5.1 that this is without loss of generality for an agent who faces an information processing constraint.
Fourth, we assume that the agent is willing to follow the principal’s recommended stopping rule whenever she is indifferent between that rule and another equally optimal stopping rule. However, it will become clear that the equilibrium strategy can be slightly modified to induce strong incentive to follow the stopping rule (see the discussion following Proposition 2).
Lastly, note that our setup says almost nothing about the agent’s motives for acquiring information. The function $\hat{u}$ describes the value the agent places on the information she receives, and the only meaningful restriction we have imposed as that it be convex (reflecting the fact that all information is at least weakly beneficial). The function $\hat{u}$ could arise from the rational anticipation of the information’s instrumental value, as described above, but could also arise from mistaken beliefs about the value of information, or non-instrumental preferences for information acquisition (e.g. a preference for the early resolution of uncertainty).
@@ 3 Optimal Policy
Throughout this section, we assume that it is prohibitively costly for the agent to perfectly learn the true state, to the point that the agent would prefer to learn nothing at all if confronted with only these two possibilities.
Assumption 1.
We assume that
$\hat{u}(\bar{q}_{0})-\frac{\bar{\kappa}}{\chi}H(\bar{q}_{0})>\sum_{x\in X}\bar{q}_{0,x}(\hat{u}(e_{x})-\frac{\bar{\kappa}}{\chi}H(e_{x})),$
where $e_{x}\in\mathcal{P}(X)$ has full support on state $x\in X$.
This assumption will imply that perfectly learning the state with probability one cannot occur in our principal-agent problem, as the principal cannot simultaneously satisfy the agent’s participation constraint and the restriction on the rate of information acquisition, (2).
3.1 A Relaxed Problem
We start by defining a relaxed version of the principal’s problem. Any probability measure and stopping rule the principal can implement will induce a probability measure over beliefs the agent will hold when she chooses to stop (i.e. a law for $q_{\tau}$). We study a version of the principal’s problem in this space. Let $\mathcal{P}(\mathcal{P}(X))$ denote the space of probability measures of beliefs on $\mathcal{P}(X)$, and, given any feasible policy chosen by the principal, let $\pi\in\mathcal{P}(\mathcal{P}(X))$ denote the probability measure of the stopped belief $q_{\tau}$. The following lemma describes the date-zero participation constraint of the agent and an upper bound on the total engagement associated with the measure $\pi$.
@@ 3.1 A Relaxed Problem
$\forall(P,I^{H},\tau)\in\mathcal{A}(q_{0})$ satisfying the agent’s optimal stopping in Definition 1, the following conditions are satisfied:
@@
    $\mathbb{E}^{\pi}[\widehat{u}(q)]-\bar{\kappa}\mathbb{E}^{P}[\tau|\mathcal{F}_{0}]\geq\widehat{u}(\bar{q}_{0})$, and
@@
Condition (1): let $\tau^{\prime}\equiv 0$, then the agent’s utility from stopping is $\widehat{\mu}(q_{0})$. The optimality of $\tau$ implies condition (i).
Condition (2): $\mathrm{d}I_{t}^{H}\leq\chi\mathrm{d}t$implies $\mathbb{E}^{P}[I_{\tau}^{H}|\mathcal{F}_{0}]=\mathbb{E}^{P}\left[\int_{0}^{\tau}\mathrm{d}I_{t}^{H}|\mathcal{F}_{0}\right]\leq\mathbb{E}^{P}\left[\int_{0}^{\tau}\chi\mathrm{d}t|\mathcal{F}_{0}\right]=\chi\mathbb{E}^{P}[\tau|\mathcal{F}_{0}]$. ∎
@@
$\mathbb{E}^{\pi}[\widehat{u}(q)-\hat{u}(\bar{q}_{0})]\geq\bar{\kappa}\mathbb{E}^{P}[\tau|\mathcal{F}_{0}]\geq\frac{\bar{\kappa}}{\chi}\mathbb{E}^{\pi}[H(q)-H(\bar{q}_{0})].$
Let us define the principal’s relaxed optimization problem as incorporating only this combined constraint:
$\bar{J}(q_{0})=\sup_{\begin{subarray}{c}\pi\in\mathcal{P}(\mathcal{P}(X)):\mathbb{E}^{\pi}[q]=q_{0}\end{subarray}}$ $\mathbb{E}^{\pi}[G(q)-G(\bar{q}_{0})]$ (R)
$s.t.$ $\frac{\bar{\kappa}}{\chi}\mathbb{E}^{\pi}[H(q)-H(\bar{q}_{0})]\leq\mathbb{E}^{\pi}[\widehat{\mu}(q)-\widehat{\mu}(\bar{q}_{0})].$
@@
Equation R has a solution $\pi^{*}$ with finite support. Either $\bar{J}(q_{0})=0$ or $\bar{J}(q_{0})>0$ and there exists $\lambda>0$ s.t. all $\pi^{*}$ satisfy
$\pi^{*}\in\arg\max_{\pi\in\mathcal{P}(\mathcal{P}(X)):\mathbb{E}^{\pi}[q]=\bar{q}_{0}}\mathbb{E}^{\pi}\left[\hat{u}(q)-\frac{\bar{\kappa}}{\chi}H(q)+\lambda G(q)\right].$
In both cases, $\frac{\bar{\kappa}}{\chi}\mathbb{E}^{\pi^{*}}[H(q)-H(\bar{q}_{0})]=\mathbb{E}^{\pi^{*}}[\widehat{\mu}(q)-\widehat{\mu}(\bar{q}_{0})]$ and $\mathbb{E}^{\pi^{*}}[\bar{J}(q)]=0$.
@@
This proposition demonstrates that $\pi^{*}$ is the solution to a static rational inattention problem, with a general UPS cost function (of the sort studied by Caplin et al. (2022)). Those authors show that a necessary condition for $\pi^{*}$ is that it concavifies the function $\widehat{\mu}-\frac{\bar{\kappa}}{\chi}H+\lambda G$ [^12] Related results appear in earlier working papers by those authors and in the Bayesian persuasion literature. Engagement is infeasible ($\bar{J}(q_{0})=0$) whenever the only $\pi$ satisfying the constraint are ones with $\mathbb{E}^{\pi}[G(q)-G(\bar{q}_{0})]=0$, which is to say that it is not possible for the principal to induce information acquisition in the dimensions he cares about. When $G$ is strictly convex (meaning that all information acquisition is at least somewhat beneficial to the principal, as in the $G=H$ case), this implies $\text{Supp}(\pi)=\{\bar{q}_{0}\}$. When $\text{Supp}(\pi)=\{\bar{q}_{0}\}$, we will say that $\pi^{*}$ is degenerate, and otherwise say that $\pi^{*}$ is non-degenerate. [^13] When $G$ is not strictly convex, meaning that there are some dimensions of potential information acquisition the principal does not value, it is possible to have $J(\bar{q}_{0})=0$ with a non-degenerate $\pi^{*}$. However, the information acquired in this case generates no surplus for either the principal or the agent.
@@ 3.2 Implementation
Take any non-degenerate $\pi\in\mathcal{P}(\mathcal{P}(X))$ s.t. $\mathbb{E}^{\pi}[q]=\bar{q}_{0}$, and define the stochastic process $q_{t}$ as:
@@
where $Q\in\mathcal{P}(X)$ is a random variable distributed according to $\pi$ and $N_{\alpha}(t)$ is an independent Poisson counting process with parameter $\alpha$. Here, $q_{t}$ is a compound Poisson process that jumps according to $\pi$ at rate $\alpha$. We call such process $q_{t}$ an $\alpha-$dilution of $\pi$ [^14] Pomatto et al. (2018) first introduce the notion of dilution. They define the dilution of an information structure $\pi$ as “producing $\pi$ with probability $\alpha$ and uninformative signal with probability $1-\alpha$.” (the same notion appeared in Bloedel and Zhong (2020)). Our notion of $\alpha$-dilution is essentially the repetition of a dilution in continuous time.
@@
For all non-degenerate $\pi\in\mathcal{P}(\mathcal{P}(X))$ that satisfy the constraint in (R), let $\alpha=\frac{\chi}{\mathbb{E}^{\pi}[H(q)-H(\bar{q}_{0})]}$ and let $q_{t}$ be the $\alpha$-dilution of $\pi$. Then $q_{t}$ is feasible in the problem of Definition 1 and implements utility level $\mathbb{E}^{\pi}[G(q)-G(\bar{q}_{0})]$.
@@
Combining Lemma 1, Proposition 1, and Proposition 2, we obtain the main characterization of the optimal policy:
@@
$\forall\bar{q}_{0}\in\mathcal{P}(X)$, there exists $\pi^{*}\in\mathcal{P}(\mathcal{P}(X))$ with finite support solving Equation R. If Supp$(\pi^{*})=\{\bar{q}_{0}\}$, the agent will immediately stop and any feasible policy is optimal. Otherwise, let $\alpha^{*}=\frac{\chi}{\mathbb{E}^{\pi^{*}}[H(q)-H(\bar{q}_{0})]}$, and let $(P^{*},\tau^{*})$ be the law and jumping time of the $\alpha^{*}$-dilution of $\pi^{*}$. Then, $(P^{*},I_{t}^{H*}=\chi t,\tau^{*})$ is an optimal solution to the principal’s problem (Definition 1).
There are generally many optimal policies in the principal’s problem. First, there may be multiple $\pi^{*}$ that solve the relaxed principal’s problem (although uniqueness in static rational inattention problems is guaranteed under certain additional assumptions). Second, there are many stochastic processes $q_{t}$ (equivalently, laws $P$) and stopping times $\tau$ that induce the same law for $q_{\tau}$; provided that this law is equal to $\pi^{*}$ and that incentive compatibility and the bounds on information acquisition are satisfied, all such policies are optimal. However, it is not the case that anything goes, as we show in section 4.
3.3 The agent-optimal benchmark and belief polarization
We will compare our results to a benchmark in which the agent and not the principal chooses the probability space and martingale belief process (subject to (2)). This benchmark is a special case of the more general models described in Hébert and Woodford (2023) and Zhong (2022).
Those authors show that the optimal policies of this benchmark dynamic model are also equivalent to the solution to a static rational inattention problem. That is, under the optimal policies given the initial belief $\bar{q}_{0}\in\mathcal{P}(X)$,
$\mathbb{E}^{P}[\hat{u}(q_{\tau})-\bar{\kappa}\tau|\mathcal{F}_{0}]$ $=V^{B}(\bar{q}_{0})$
$=\max_{\pi\in\mathcal{P}(\mathcal{P}(X)):\medspace E^{\pi}[q]=\bar{q}_{0}}\mathbb{E}^{\pi}[\hat{u}(q)-\frac{\bar{\kappa}}{\chi}(H(q)-H(\bar{q}_{0}))].$

### v6

@@ 2.3 Discussion
The generalized entropy functions $G$ and $H$ and value of information $\hat{u}$ play conceptually distinct roles in our model. The function $H$ governs the agent’s information processing constraint. It determines which kinds of information are relatively easy or difficult for the agent to process. The function $\hat{u}$ describes the information the agent would like to receive, and the function $G$ describes the information the principal would like the agent to receive.
To simplify the discussion that follows, we will assume that the state space has a product structure, $X=X_{1}\times X_{2}$, and that the agent values only information about $x_{1}\in X_{1}$ (i..e $\hat{u}$ depends only on the marginal posterior over $X_{1}$). The agent benefits from learning about $x_{2}\in X_{2}$ only if doing so makes it easier to learn about $x_{1}\in X_{1}$. Whether or not this is the case is governed by the $H$ function, and our general framework admits both possibilities. [^10] The specific property of the $H$ function that governs whether or not learning about $x_{2}$ makes it cheaper to learn about $x_{1}$ is what hebert2023information call “R-monotonicity.” We interpret $x_{1}$ as the useful content and $x_{2}$ as content that has no intrinsic value to the user. Throughout the paper, we assume that the user “passively” learns about both $x_{1}$ and $x_{2}$, as long as the principal supplies both types of content. In the internet platform context, this can be interpreted as reflecting “stealth marketing” or “click-baiting” behavior that makes ads indistinguishable from relevant content. However, even if the ads are obvious and the user can effortlessly avoid them, she still finds it optimal to fully process information about $x_{2}$ in equilibrium. [^11] This statement has been formally proved in the working paper version of the current paper (see Proposition 8 of hébert2025engagementmaximization).
We focus our discussion on the extent to which $G$ differs from $H$ and $\hat{u}$. One leading case of interest is when G=HG=H. [^12] The case in which $G$ is proportional to $\hat{u}$ is essentially identical to the $G=H$ case, and for this reason we will discuss only the latter. We interpret this case as reflecting a kind of indifference on the part of the principal: he does not care what the agent learns about, only that she learn as much as possible. In equilibrium, the constraint (1) will bind; as a result, “learning as much as possible” is equivalent to “spending as much time learning as possible.” [^13] The authors would like to thank Emir Kamenica for suggesting this interpretation. Note that although time spent and information processed are equivalent when (1) binds, they are not equivalent more generally (e.g. if the principal sends no signals to waste time), which is why we describe the equivalence as “time spent learning.” We will show in this case that the signals sent by the principal maximize the utility of a hypothetical agent with a lower cost of delay.
A second leading case is one in which $G$ measures only information acquisition that is irrelevant to the agent’s decision problem (i.e. it depends only on the marginal posterior of $X_{2}$). We interpret this case as capturing a multi-dimensional conflict between the principal and agent over both how much the agent should learn and what the agent should learn about. In this case, the signals sent by the principal are not equivalent to those that would be chosen by a hypothetical agent with a lower cost of delay; the principal also distorts the signals so that they are more informative about $X_{2}$. In the special case in which $H$ is proportional to Shannon’s entropy, we derive a sharp result: the principal provides the agent with information about $x_{1}\in X_{1}$ that is identical to what the agent would optimally choose for herself, plus additional information about $x_{2}\in X_{2}$ that the agent would never choose to receive.
There are also intermediate situations of interest. The information the principal would like the agent to acquire might be of some use to the agent without being neutral or aligned with what the agent would like to learn (i.e. $G$ can depend in part of the posterior over $X_{1}$ without being a linear combination of $\hat{u}$ and $H$). Our two leading cases generate sharp predictions precisely because they are extreme.
@@ 3 Optimal Policy
We start by defining a relaxed version of the principal’s problem. Any probability measure and stopping rule the principal can implement will induce a probability measure over beliefs the agent will hold when she chooses to stop (i.e. a law for $q_{\tau}$). Define $\Pi(\bar{q_{0}})=\{\pi\in\mathcal{P}(\mathcal{P}(X)):\mathbb{E}^{\pi}[q]=\bar{q}_{0}\}$ as the set of probability measures over posterior beliefs that are consistent with the initial prior. Under any feasible policy $(P,\tau)$ chosen by the principal, the law of the stopped belief $q_{\tau}$ induced by this policy will be an element of $\Pi(\bar{q}_{0}$), by the martingale property of beliefs.
The following lemma describes the date-zero participation constraint of the agent (i.e., incentive compatibility w.r.t. stopping at time 0) and an upper bound on the total engagement.
@@
$\forall(P,\tau)\in\mathcal{A}(\bar{q}_{0})$ satisfying the agent’s optimal stopping constraint in Definition 1, the law $\pi$ of the stopped belief $q_{\tau}$ satisfies:
@@
    $\mathbb{E}^{\pi}[\widehat{u}(q)]\geq\kappa\mathbb{E}^{P}[\tau|\mathcal{F}_{0}]+\widehat{u}(\bar{q}_{0})$, and
@@
See Appendix A.1. ∎
@@
$\mathbb{E}^{\pi}[\widehat{u}(q)-\widehat{u}(\bar{q}_{0})]\geq\kappa\mathbb{E}^{P}[\tau|\mathcal{F}_{0}]\geq\frac{\kappa}{\chi}\mathbb{E}^{\pi}[H(q)-H(\bar{q}_{0})].$
Let us define the principal’s relaxed optimization problem as a maximization over $\Pi(\bar{q}_{0})$ incorporating only this combined constraint:
$\bar{J}(\bar{q}_{0})$ $=\sup_{\pi\in\Pi(\bar{q}_{0})}\mathbb{E}^{\pi}[G(q)-G(\bar{q}_{0})]$ (R)
$s.t.\$$\frac{\kappa}{\chi}\mathbb{E}^{\pi}[H(q)-H(\bar{q}_{0})]\leq\mathbb{E}^{\pi}[\widehat{u}(q)-\widehat{u}(\bar{q}_{0})].$
@@
We assume that (1) it is possible for the principal to benefit from engagement and (2) it is prohibitively costly for the agent to learn enough information to achieve the principal’s first best, to the point that the agent would prefer to learn nothing at all if confronted with only these two possibilities.
Assumption 1.
The principal can benefit from engagement: $\max_{\pi^{\prime}\in\Pi(\bar{q}_{0})}\mathbb{E}^{\pi^{\prime}}[G(q)]>G(\bar{q}_{0})$. However, it is not possible for the principal to achieve the principal’s first best: if $\pi\in\arg\max_{\pi^{\prime}\in\Pi(\bar{q}_{0})}\mathbb{E}^{\pi^{\prime}}[G(q)]$, then
$\widehat{u}(\bar{q}_{0})-\frac{\kappa}{\chi}H(\bar{q}_{0})>\mathbb{E}^{\pi}\left[\widehat{u}(q)-\frac{\kappa}{\chi}H(q)\right].$
This rules out the possibility that the principal either has no desire for the agent to acquire information or the possibility of satiating this desire. It immediately implies that the constraint in (R) will bind.
Define the function $\mathcal{L}$ as the Lagrangian associated with the dual of (R),
$\mathcal{L}(\pi,\lambda)=\mathbb{E}^{\pi}\left[\widehat{u}(q)-\frac{\kappa}{\chi}H(q)+\lambda G(q)\right].$
The following proposition characterizes the solutions of (R).
@@
A solution to (R) exists. Additionally,
  • there exists a $\lambda\geq 0$ such that if $\pi^{*}$ is a solution to (R) then
    $\pi^{*}\in\arg\max_{\pi\in\Pi(\bar{q}_{0})}\mathcal{L}(\pi,\lambda),$
  • if $\pi^{*}$ is a solution to (R), then the constraint binds,
    $\frac{\kappa}{\chi}\mathbb{E}^{\pi^{*}}[H(q)-H(\bar{q}_{0})]=\mathbb{E}^{\pi^{*}}[\widehat{u}(q)-\widehat{u}(\bar{q}_{0})],$
  • if $\pi^{*}$ is a solution to (R), then starting (R) from any belief in the support of $\pi^{*}$, it is not possible to induce engagement, $\mathbb{E}^{\pi^{*}}[\bar{J}(q)]=0$, and
  • at least one solution to (R) has finite support.
@@
This proposition demonstrates that $\pi^{*}$ is the solution to a static rational inattention problem, with a general UPS cost function (of the sort studied by caplin2017shannon). Those authors show that a necessary condition for $\pi^{*}$ is that it concavifies the function $\widehat{u}-\frac{\kappa}{\chi}H+\lambda G$. [^14] Related results appear in earlier working papers by those authors and in the Bayesian persuasion literature. Here, $\hat{u}$ captures the benefit of information to the agent, $\frac{\kappa}{\chi}H$ captures the costly delay required to acquire the information, and $\lambda G$ captures the benefit of the information acquisition to the principal, converted to the units of the agent’s utility at the rate $\lambda$.
Engagement is infeasible ($\bar{J}(\bar{q}_{0})=0$) whenever the only $\pi$ satisfying the constraint are ones with $\mathbb{E}^{\pi}[G(q)-G(\bar{q}_{0})]=0$, which is to say that it is not possible for the principal to induce information acquisition in the dimensions he cares about. If $G$ is strictly convex (meaning that all information acquisition is at least somewhat beneficial to the principal, as in the $G=H$ case), this implies $\text{Supp}(\pi)=\{\bar{q}_{0}\}$. When $\text{Supp}(\pi)=\{\bar{q}_{0}\}$, we will say that $\pi^{*}$ is degenerate, and otherwise say that $\pi^{*}$ is non-degenerate. [^15] When $G$ is not strictly convex, meaning that there are some dimensions of potential information acquisition the principal does not value, it is possible to have $J(\bar{q}_{0})=0$ with a non-degenerate $\pi^{*}$. However, the information acquired in this case generates no surplus for either the principal or the agent.
@@ 3.1 Implementation
Take any non-degenerate $\pi\in\Pi(\bar{q}_{0})$, and define the stochastic process $q_{t}$ as:
@@
where $Q\in\mathcal{P}(X)$ is a random variable distributed according to $\pi$ and $N_{\alpha}(t)$ is an independent Poisson counting process with parameter $\alpha$. Here, $q_{t}$ is a compound Poisson process that jumps according to $\pi$ at rate $\alpha$. We call such process $q_{t}$ an $\alpha$-dilution of $\pi$.161616pomatto2018cost first introduce the notion of dilution. They define the dilution of an information structure $\pi$ as “producing $\pi$ with probability $\alpha$ and uninformative signal with probability $1-\alpha$.” (the same notion appeared in bloedel2020cost). Our notion of $\alpha$-dilution is essentially the repetition of a dilution in continuous time.
@@
For all non-degenerate $\pi\in\Pi(\bar{q}_{0})$ that satisfy the constraint in (R), let $\alpha=\frac{\chi}{\mathbb{E}^{\pi}[H(q)-H(\bar{q}_{0})]}$ and let $q_{t}$ be the $\alpha$-dilution of $\pi$. Let $P$ be the law of $q_{t}$ and let $\tau=\inf\{t>0:q_{t}\neq\bar{q}_{0}\}$. Then $(P,\tau)$ is feasible in (P) and implements utility level $\mathbb{E}^{\pi}[G(q)-G(\bar{q}_{0})]$.
@@
Note that Proposition 2 immediately implies that $J=\bar{J}$; hence, we will not distinguish between the two functions. Combining Lemma 1, Proposition 1, and Proposition 2, we obtain the main characterization of the optimal policy:
@@
[…]
There are generally many optimal policies in the principal’s problem. First, there may be multiple $\pi^{*}$ that solve the relaxed principal’s problem (although uniqueness in static rational inattention problems is guaranteed under certain additional assumptions). Second, there are many stochastic processes $q_{t}$ (equivalently, laws $P$) and stopping times $\tau$ that induce the same law for $q_{\tau}$; provided that this law is equal to $\pi^{*}$ and that incentive compatibility and the bounds on information acquisition are satisfied, all such policies are optimal. However, it is not the case that anything goes, as we show in section˜4.
$\mathbb{E}^{P^{*}}[\widehat{u}(q_{\tau^{*}})-\kappa\tau^{*}|\mathcal{F}_{0}]=\widehat{u}(\bar{q}_{0}).$
Strikingly, the agent is no better off receiving information from an engagement-maximizing principal than if she could not receive any information at all. The principal extracts the full surplus generated by the ability to produce information and apply it in the agent’s decision problem, despite the agent’s ability to choose when to stop and act. The principal, in maximizing the engagement of the agent, minimizes the agent’s welfare subject to a participation constraint. Note, however, that the principal’s ability to extract all of the surplus depends critically on the possibility of jumps in the belief process. We provide an example in Proposition 7 in which beliefs are required to be continuous and the agent-optimal outcome is attained.
3.2 The agent-optimal benchmark and extreme beliefs
We will compare our results to a benchmark in which the agent and not the principal chooses the probability space and martingale belief process (subject to (1)). This benchmark is a special case of the more general models described in hebert2019rational and zhong2017jumps.
Those authors show that the optimal policies of this benchmark dynamic model are also equivalent to the solution to a static rational inattention problem. That is, under the agent-optimal policies $(P_{A},\tau_{A})$, given the initial belief $\bar{q}_{0}\in\mathcal{P}(X)$,
$\mathbb{E}^{P_{A}}[\widehat{u}(q_{\tau_{A}})-\kappa\tau_{A}|\mathcal{F}_{0}]$ $=V^{B}(\bar{q}_{0})$
$=\max_{\pi\in\Pi(\bar{q}_{0})}\mathbb{E}^{\pi}[\widehat{u}(q)-\frac{\kappa}{\chi}(H(q)-H(\bar{q}_{0}))].$